# Aligning Visual Foundation Encoders to Tokenizers for Diffusion Models

### v1

@@ Abstract
In this work, we propose aligning pretrained visual encoders to serve as tokenizers for latent diffusion models in image generation. Unlike training a variational autoencoder (VAE) from scratch, which primarily emphasizes low-level details, our approach leverages the rich semantic structure of foundation encoders. We introduce a three-stage alignment strategy: (1) freeze the encoder and train an adapter and a decoder to establish a semantic latent space; (2) jointly optimize all components with an additional semantic preservation loss, enabling the encoder to capture perceptual details while retaining high-level semantics; and (3) refine the decoder for improved reconstruction quality. This alignment yields semantically rich image tokenizers that benefit diffusion models. On ImageNet $\times256$, our tokenizer accelerates the convergence of diffusion models, reaching a gFID of 1.90 within just 64 epochs, and improves generation both with and without classifier-free guidance. Scaling to LAION, a 2B-parameter text-to-image model trained with our tokenizer consistently outperforms FLUX VAE under the same training steps. Overall, our method is simple, scalable, and establishes a semantically grounded paradigm for continuous tokenizer design. Project Page.
@@ 1 Introduction
We demonstrate the effectiveness of our method on both the ImageNet $\times256$ dataset (Deng et al., 2009) and the LAION dataset (Schuhmann et al., 2022) by training diffusion models using our continuous tokenizers. On ImageNet, our semantic tokenizer accelerates the convergence of diffusion models and improves generation quality over previous methods, both with and without classifier-free guidance (CFG), as well as in unconditional generation settings. To further validate scalability, we train a 2B-parameter text-to-image diffusion model on LAION and show that it converges significantly faster than the FLUX VAE (Labs, 2024). We conjecture that these benefits arise from a well-grounded semantic latent space, which provides a more structured representation.
@@ 2.2 Image Tokenizers for Generative Models
Image tokenizers are essential for scaling visual generation. They use an encoder–decoder framework to map images into compact latent spaces where generative models can operate more effectively and efficiently (Rombach et al., 2022). A common practice in training tokenizers is to rely on reconstruction losses such as L1 loss, perceptual loss, and adversarial loss (Rombach et al., 2022; Kouzelis et al., 2025; Wen et al., 2025; Wu et al., 2025; Yu et al., 2024; Skorokhodov et al., 2025; Hansen-Estruch et al., 2025; Silvestri & Ambrogioni, 2025; Van Den Oord et al., 2017). Depending on the design, tokenizers are either continuous, where the latent distribution is regularized by a KL loss, or discrete, where quantization with a codebook is enforced via a VQ loss. While these methods achieve faithful reconstruction fidelity, the resulting latent space is often dominated by fine-grained details because they mainly rely on reconstruction loss, which can hinder generative performance.
@@
Pretrained Encoders as Discrete Tokenizers. The use of pretrained encoders in tokenizers has also been studied, but primarily in the context of discrete tokenizers for autoregressive models. One line of work treats the pretrained encoder as a fixed feature extractor without fine-tuning it to encode perceptual details (Zheng et al., 2025; Chen et al., 2025c; Xie et al., 2024; Wang et al., 2024). Another line of work fine-tunes pretrained encoders with additional architectural designs, such as introducing extra encoders (Qu et al., 2025; Jiao et al., 2025) or decoders (Han et al., 2025), or splitting encoder features into separate vocabularies (Song et al., 2025). A third direction leverages contrastive learning, fine-tuning the pretrained encoder with image–text supervision to capture semantic alignment (Wu et al., 2024; Ma et al., 2025; Zhao et al., 2025; Lin et al., 2025). However, this strategy primarily assumes that the encoder is a language-aligned model (e.g., SigLIP 2).
@@ 4.1 Ablation Study
We also test a variant Applied Pre-Adapter where the semantic preservation loss is applied directly on the outputs of the pretrained encoder (before adapter). This approach partially preserves semantics, and the generation quality lags behind Stage 1 + Stage 2 variant. This may be because leaving the adapter unregularized gives it too much flexibility, leading to a loss of semantic structure.
@@ 4.2.1 Class-Conditional Generation
Convergence Speed. We compare the convergence speed of generative model training using our tokenizer against VA-VAE. As shown in Fig. 4 right, our approach consistently achieves better gFID across training iterations. By aligning with a pretrained encoder rather than relying on regularization, our method provides a more semantically structured latent space. This leads to roughly 5x faster training, requiring only $\sim$60K steps compared to VA-VAE’s 300K steps for comparable quality, when evaluated with 30 sampling steps. See Appendix for qualitative comparisons (Sec. B.3).
@@ 4.2.2 Unconditional Generation
Table 4: System-Level Comparison. We compare with VAR (Tian et al., 2024), MagViT-v2 (Yu et al., 2023), MAR (Li et al., 2024), l-DeTok (Yang et al., 2025), MaskDiT (Zheng et al., 2023), DiT (Peebles & Xie, 2023), SiT (Ma et al., 2024), FasterDiT (Yao et al., 2024), MDT (Gao et al., 2023a), MDTv2 (Gao et al., 2023b), REPA (Yu et al., 2025), CausalFusion (Deng et al., 2024), MAETok (Chen et al., 2025a), and VA-VAE (Yao et al., 2025). Gray and purple regions refer to LightningDiT trained for 64 epochs (80K training steps, no QKNorm) and 800 (1M training steps, with QKNorm) epochs, respectively. Bold numbers indicate the best result in each color block.
@@
LightningDiT Ours $256\times 32$ 0.26 800 2.04 206.2 0.76 0.67 1.37 293.6 0.79 0.65
@@
LightningDiT Ours $256\times 32$ 0.26 64 3.71 148.9 0.77 0.62 1.90 260.9 0.81 0.61
@@ 4.2.3 Reconstruction
As shown in Tab. 3, our method achieves competitive reconstruction with 32 channels but lags behind VA-VAE (CNN encoder) at 64 channels. Lowering the semantic preservation loss weight and increasing learning rate in stage 2 improves reconstruction to match VA-VAE, with only a slight drop in generation performance (still surpassing VA-VAE). See Appendix (Sec. B.2) for details.
@@ 4.4 Scale-Up Experiments on Text-to-Image Generation
For comparison, Fig. 5 presents results from generative models trained with our tokenizer and with FLUX VAE for 100K steps at $\times256$ resolution. Our method produces images with better coherence, stronger text alignment, and competitive visual quality. Quantitative results (Tab. 5) further confirm the advantage of our approach. Fig. 6 presents our generated results at 512 resolution. Notably, our tokenizer is trained only on 256-resolution images, demonstrating its ability to generalize to unseen resolutions. These results suggest that our approach scales effectively and can potentially serve as a strong alternative to FLUX VAE in large-scale training. See Appendix (Sec. B.3-B.5) for additional quantitative results on two more prompt sets, as well as qualitative results, including convergence speed comparisons and generated images across different resolutions and aspect ratios.
Table 5: Quantitative Comparison on Text-to-Image (T2I) Generation with FLUX VAE. We compare on COCO Prompt 6K, which has 6K captions sampled from the COCO validation set. Each T2I model is trained for 100K steps and evaluated at $\times256$ resolution. The rFID is computed using 200K randomly sampled images from the COYO-700M dataset (Minwoo et al., 2022).
@@ 5 Limitations and Discussions
6 Acknowledgement
This work was done while Bowei Chen and Tianyuan Zhang were interning at Adobe. We would like to thank Hailin Jin for his valuable support throughout the internship.
@@ Appendix A Implementation Details
Appendix B More Experiments
B.1 Other Pretrained Encoders
@@
B.2 Reconstruction
@@
B.3 Convergence Speed
Figs.  9, Fig. 10, and Fig. 11 present additional comparisons between our method and VA-VAE on the ImageNet $\times256$ dataset. Our method converges faster, indicating that aligning and preserving semantics in the pretrained encoder is more effective than semantic regularization.
@@ B.3 Convergence Speed
B.4 More Quantitative Results
@@
Tab. 8 reports quantitative results on two additional prompt sets, showing that our tokenizer consistently outperforms FLUX VAE in generation performance.
B.5 More Qualitative Results

### v2

# AlignTok: Aligning Visual Foundation Encoders to Tokenizers for Diffusion Models

@@ Abstract
In this work, we propose aligning pretrained visual encoders to serve as tokenizers for latent diffusion models in image generation. Unlike training a variational autoencoder (VAE) from scratch, which primarily emphasizes low-level details, our approach leverages the rich semantic structure of foundation encoders. We introduce a three-stage alignment strategy called AlignTok: (1) freeze the encoder and train an adapter and a decoder to establish a semantic latent space; (2) jointly optimize all components with an additional semantic preservation loss, enabling the encoder to capture perceptual details while retaining high-level semantics; and (3) refine the decoder for improved reconstruction quality. This alignment yields semantically rich image tokenizers that benefit diffusion models. On ImageNet $\times256$, our tokenizer accelerates the convergence of diffusion models, reaching a gFID of 1.90 within just 64 epochs, and improves generation both with and without classifier-free guidance. Scaling to LAION, text-to-image models trained with our tokenizer consistently outperforms FLUX VAE and VA-VAE under the same training steps. Overall, our method is simple, scalable, and establishes a semantically grounded paradigm for continuous tokenizer design.
@@ 1 Introduction
We demonstrate the effectiveness of our method on both the ImageNet $\times256$ dataset (Deng et al., 2009) and the LAION dataset (Schuhmann et al., 2022) by training diffusion models using our continuous tokenizers. On ImageNet, our semantic tokenizer accelerates the convergence of diffusion models and improves generation quality over previous methods, both with and without classifier-free guidance (CFG), as well as in unconditional generation settings. To further validate scalability, we train text-to-image diffusion models on LAION and show that they converge significantly faster than the FLUX VAE (Labs, 2024) and VA-VAE (Yao et al., 2025). We conjecture that these benefits arise from a well-grounded semantic latent space, which provides a more structured representation.
@@ 2.2 Image Tokenizers for Generative Models
Image tokenizers are essential for scaling visual generation. They use an encoder–decoder framework to map images into compact latent spaces where generative models can operate more effectively and efficiently (Rombach et al., 2022). A common practice in training tokenizers is to rely on reconstruction losses such as L1 loss, perceptual loss, and adversarial loss (Rombach et al., 2022; Kouzelis et al., 2025; Wen et al., 2025; Wu et al., 2025; Yu et al., 2024; Skorokhodov et al., 2025; Hansen-Estruch et al., 2025; Silvestri and Ambrogioni, 2025; Van Den Oord et al., 2017). Depending on the design, tokenizers are either continuous, where the latent distribution is regularized by a KL loss, or discrete, where quantization with a codebook is enforced via a VQ loss. While these methods achieve faithful reconstruction fidelity, the resulting latent space is often dominated by fine-grained details because they mainly rely on reconstruction loss, which can hinder generative performance.
@@
Pretrained Encoders as Discrete Tokenizers. The use of pretrained encoders in tokenizers has also been studied, but primarily in the context of discrete tokenizers for autoregressive models. One line of work treats the pretrained encoder as a fixed feature extractor without fine-tuning it to encode perceptual details (Zheng et al., 2025a; Chen et al., 2025c; Xie et al., 2024; Wang et al., 2024). Another line of work fine-tunes pretrained encoders with additional architectural designs, such as introducing extra encoders (Qu et al., 2025; Jiao et al., 2025) or decoders (Han et al., 2025), or splitting encoder features into separate vocabularies (Song et al., 2025). A third direction leverages contrastive learning, fine-tuning the pretrained encoder with image–text supervision to capture semantic alignment (Wu et al., 2024; Ma et al., 2025; Zhao et al., 2025; Lin et al., 2025). However, this strategy primarily assumes that the encoder is a language-aligned model (e.g., SigLIP 2).
@@
Relationship to RAE. RAE Zheng et al. (2025b) proposes to directly adopt a pretrained foundation encoder, without fine-tuning, as the tokenizer encoder. While this preserves strong semantic representations, it results in a high-dimensional latent space with a large channel dimension. Diffusion models are known to struggle in such high-channel latent spaces, as optimization becomes unstable and noise scheduling becomes less effective. To mitigate this issue, RAE introduces techniques such as dimension-dependent noise shifting and a widened diffusion head to stabilize training in high-dimensional latents. Although RAE demonstrates strong performance for downstream generative model training, its reconstruction quality is limited by the frozen encoder, which is not optimized for faithful pixel-level reconstruction. In contrast, our approach fine-tunes the encoder, significantly improving reconstruction fidelity while maintaining semantic alignment. We believe that combining RAE’s high-channel semantic latent space with our alignment and fine-tuning strategy could offer the best of both worlds: strong semantic representations together with high-quality reconstruction.
Relationship to REPA-E. While both REPA-E Leng et al. (2025) and our method leverage a vision foundation model, it is not straightforward to claim an overall advantage of one approach over the other because the two methods follow different setups and can in fact be combined. Specifically, REPA-E trains the tokenizer and the diffusion model jointly in an end-to-end manner. Even under this end-to-end regime, the authors show that initializing the tokenizer with an existing tokenizer works better than random initialization; specifically, their best model (E2E-VAE) initializes the tokenizer from VA-VAE. In contrast, our method focuses solely on training the tokenizer itself. This makes our tokenizer a drop-in replacement for the VA-VAE component inside REPA-E: since REPA-E already benefits from initializing its tokenizer with VA-VAE, it could similarly initialize with our tokenizer as a potentially stronger starting point for its end-to-end training.
@@ 4.1 Ablation Study
We test a variant Applied Pre-Adapter where the semantic preservation loss is applied directly on the outputs of the pretrained encoder (before adapter). This approach partially preserves semantics, and the generation quality lags behind Stage 1 + Stage 2 variant. This may be because leaving the adapter unregularized gives it too much flexibility, leading to a loss of semantic structure. In addition, we also test a variant Cosine Loss that replaces the L2 loss with a cosine similarity loss, which applies the semantic constraint with explicit normalization. This variant achieves performance comparable to L2 loss. This suggests that, at least in our current setup, both losses lead to similar behavior, indicating that the performance differences are not dominated by the specific choice of L2 versus cosine loss.
@@ 4.2.1 Class-Conditional Generation
Convergence Speed. We compare the convergence speed of generative model training using our tokenizer against VA-VAE. As shown in Fig. 4 right, our approach consistently achieves better gFID across training iterations. By aligning with a pretrained encoder rather than relying on regularization, our method provides a more semantically structured latent space. This leads to roughly 5x faster training, requiring only $\sim$60K steps compared to VA-VAE’s 300K steps for comparable quality, when evaluated with 30 sampling steps. See Appendix for qualitative comparisons (Sec. C.3).
@@ 4.2.2 Unconditional Generation
Table 4: System-Level Comparison. We compare with VAR (Tian et al., 2024), MagViT-v2 (Yu et al., 2023), MAR (Li et al., 2024), l-DeTok (Yang et al., 2025), MaskDiT (Zheng et al., 2023), DiT (Peebles and Xie, 2023), SiT (Ma et al., 2024), FasterDiT (Yao et al., 2024), MDT (Gao et al., 2023a), MDTv2 (Gao et al., 2023b), REPA (Yu et al., 2025), CausalFusion (Deng et al., 2024), MAETok (Chen et al., 2025a), and VA-VAE (Yao et al., 2025). Gray and purple regions refer to LightningDiT trained for 64 epochs (80K training steps, no QKNorm) and 800 (1M training steps, with QKNorm) epochs, respectively. Bold numbers indicate the best result in each color block.
@@
LightningDiT AlignTok $256\times 32$ 0.26 800 2.04 206.2 0.76 0.67 1.37 293.6 0.79 0.65
@@
LightningDiT AlignTok $256\times 32$ 0.26 64 3.71 148.9 0.77 0.62 1.90 260.9 0.81 0.61
@@ 4.2.3 Reconstruction
As shown in Tab. 3, our method achieves competitive reconstruction with 32 channels but lags behind VA-VAE (CNN encoder) at 64 channels. Lowering the semantic preservation loss weight and increasing learning rate in stage 2 improves reconstruction to match VA-VAE, with only a slight drop in generation performance (still surpassing VA-VAE). See Appendix (Sec. C.2) for details.
@@ 4.4 Scale-Up Experiments on Text-to-Image Generation
For comparison, Fig. 5 presents results from generative models trained with our tokenizer and with FLUX VAE for 100K steps at $\times256$ resolution. Our method produces images with better coherence, stronger text alignment, and competitive visual quality. Quantitative results (Tab. 5) further confirm the advantage of our approach. Fig. 6 presents our generated results at 512 resolution. Notably, our tokenizer is trained only on 256-resolution images, demonstrating its ability to generalize to unseen resolutions. These results suggest that our approach scales effectively and can potentially serve as a strong alternative to FLUX VAE in large-scale training. See Appendix (Sec. C.3-C.5) for additional quantitative results on two more prompt sets, as well as qualitative results, including convergence speed comparisons and generated images across different resolutions and aspect ratios.
Table 5: Quantitative Comparison on Text-to-Image (T2I) Generation with FLUX VAE. Compared on COCO Prompt 6K, which has 6K captions sampled from the COCO validation set. Each 2B-parameter T2I model is trained for 100K steps and evaluated at 256×256 resolution with CFG. rFID is computed using 200K randomly sampled images from the COYO-700M dataset (Minwoo et al., 2022).
@@ 5 Limitations and Discussions
Acknowledgements. This work was done while Bowei Chen and Tianyuan Zhang were interning at Adobe. We would like to thank Hailin Jin and Jingwan (Cynthia) Lu for their valuable support throughout the internship.
Reproducibility Statement. We provide the implementation details and hyperparameters in Sec. A of the Appendix, which are sufficient to reproduce the results of our method.
@@ Appendix A Implementation Details
Linear Probing. We follow the standard ImageNet protocol widely used to assess the semantic quality of latent representations. Specifically, we freeze both the VAE encoder and the adapter, and train a single linear classifier on top of the adapter’s output features to predict ImageNet-1K classes. We use linear probing because it directly measures how much semantic information is linearly accessible in the latent space, making it a widely adopted proxy for semantic quality. The classifier is trained for 3 epochs using Adam optimizer with a learning rate of 0.001 and a batch size of 256 on the ImageNet-1K training set, and evaluated on the ImageNet validation set.
@@
Appendix B Training and Inference Cost
We show cost comparison of training in Tab. 6. Our main observations are as follows:
(1) Total GPU memory. The vanilla VAE consumes the least memory because it does not load any pretrained encoder. Our Stage 1 and Stage 3 use less memory than VA-VAE since our encoder is frozen during these stages, whereas VA-VAE loads CNN encoder (trainable) + DINOv2 encoder (frozen). Our Stage 2 consumes the most memory because it fine-tunes the DINOv2 encoder.
(2) Time per training step. The pattern mirrors memory usage for the same reason: VA-VAE incurs higher cost than our Stage 1 and 3, and our Stage 2 incurs the highest cost when the encoder is trainable.
(3) Total A100 GPU hours. The cumulative GPU training hours of our method (107.2 + 377.4 + 91.55 = 576.15) are lower than those of both Vanilla VAE and VA-VAE, demonstrating the overall training efficiency of our pipeline.
We also show cost comparison of inference in Tab. 7 and Tab. 8. Our main observations are as follows:
(1) Peak GPU Memory. At lower resolutions and smaller batch sizes, our encoder consumes more peak GPU memory than VA-VAE (ours: 1.372 GB, VA-VAE: 0.838 GB at 256 resolution with a batch size of 4). However, at higher resolutions and larger batch sizes, our encoder uses substantially less memory (ours: 3.015 GB, VA-VAE: 22.78 GB at 1024 resolution with a batch size of 8).
(2) Latency and GFLOPS. Overall, our tokenizer has higher encoding latency and GFLOPS than VA-VAE. The only exception is at 512 resolution, where our tokenizer achieves lower encoding latency.
Appendix C More Experiments
C.1 Other Pretrained Encoders
@@
Table 6: Training Cost Comparison. All measurements are obtained at an input resolution of $\times256$ on 8 NVIDIA A100 GPUs (batch size 8 per GPU). Values marked with * denote GPU-hour estimates computed from the training epochs reported in the official implementation of (Yao et al., 2025).
Tokenizer Enc. #Param Trainable Params Frozen Params Total GPU Memory Time Per Training Step Total A100 GPU Hours
Vanilla VAE 28.41 M 72.60 M 14.71 M 148.2 GB 0.360 s 800.0
VA-VAE 72.63 M 319.0 M 203.0 GB 0.534 s 2966
Our Stage 1 304.4 M 44.25 M 319.0 M 157.5 GB 0.402 s 107.2
Our Stage 2 348.6 M 319.0 M 238.2 GB 0.772 s 377.4
Our Stage 3 44.18 M 319.1 M 159.0 GB 0.412 s 91.55
Table 7: Inference Memory Consumption Comparison. We evaluate the peak GPU memory consumption under different image resolutions and batch sizes.
Tokenizer Image Resolution Encoder #Params Encoder Peak GPU Memory (GB) Decoder #Params Decoder Peak GPU Memory (GB)
batch size=1 batch size=4 batch size = 8 batch size=1 batch size=4 batch size = 8
VA-VAE 256 $\times$ 256 28.41 M 0.301 0.838 1.543 41.42 M 0.310 0.713 1.250
Ours 256 $\times$ 256 304.4 M 1.363 1.372 1.385
VA-VAE 512 $\times$ 512 28.41 M 0.841 2.959 5.794 41.42 M 0.713 2.324 4.472
Ours 512 $\times$ 512 304.4 M 1.284 1.450 1.674
VA-VAE 1024 $\times$ 1024 28.41 M 2.958 11.45 22.78 41.42 M 2.324 8.768 17.36
Ours 1024 $\times$ 1024 304.4 M 1.450 2.121 3.015
Table 8: Inference Compute Cost Comparison. We measure the GFLOPS and Latency under different image resolutions, using a single NVIDIA A100 GPU with a batch size of 1.
Tokenizer Image Resolution Encoder Decoder
#Params GFLOPS Latency #Params GFLOPS Latency
VA-VAE 256 $\times$ 256 28.41 M 69.21 6.548 ms 41.42 M 126.5 13.91 ms
Ours 256 $\times$ 256 304.4 M 77.85 13.42 ms
VA-VAE 512 $\times$ 512 28.41 M 279.2 19.44 ms 41.42 M 509.3 42.31 ms
Ours 512 $\times$ 512 304.4 M 310.4 18.92 ms
VA-VAE 1024 $\times$ 1024 28.41 M 1155 73.92 ms 41.42 M 2089 173.9 ms
Ours 1024 $\times$ 1024 304.4 M 1241 121.2 ms
C.2 Reconstruction
@@
C.3 Convergence Speed
Figs.  10, Fig. 11, and Fig. 12 present additional comparisons between our method and VA-VAE on the ImageNet $\times256$ dataset. Our method converges faster, indicating that aligning and preserving semantics in the pretrained encoder is more effective than semantic regularization.
@@ C.3 Convergence Speed
C.4 More Quantitative Results
@@
Tab. 11 reports quantitative results on two additional prompt sets. Tab. 13 and Tab. 15 present comparisons against FLUX VAE and VA-VAE on GenEval Ghosh et al. (2023). Tab. 12 and Tab. 14 further compare the main metrics under settings with and without CFG. Across all configurations, our tokenizer consistently outperforms both FLUX VAE and VA-VAE in generation quality, regardless of whether CFG is applied.
Table 12: Quantitative Comparison on Text-to-Image (T2I) Generation with FLUX VAE (Without CFG). We report metrics on diverse prompt sets for 2B-parameter T2I models trained with our tokenizer and the FLUX VAE. Each T2I model is trained for 100K steps and evaluated at 256 $\times$ 256 resolution. Our tokenizer is trained on LAION dataset. The rFID is computed using 200K randomly sampled images from the COYO-700M dataset (Minwoo et al., 2022).
Coco Prompt 6K (Lin et al., 2014) (without CFG)
FLUX VAE 0.102 48.76 0.025 0.183 0.345 -1.127 4.205 27.05 0.534
Ours 0.443 33.46 0.014 0.210 0.655 -0.581 4.963 28.87 0.658
Parti Prompt (Yu et al., 2022) (without CFG)
FLUX VAE 0.102 - - 0.186 0.378 -1.176 4.003 26.85 0.538
Ours 0.443 - - 0.206 0.622 -0.654 4.751 28.59 0.618
HPSv2 Prompt (Wu et al., 2023) (without CFG)
FLUX VAE 0.102 - - 0.170 0.372 –1.206 4.142 26.59 0.567
Ours 0.443 - - 0.192 0.628 -0.764 4.955 28.20 0.645
Table 13: GenEval Quanlitative Comparison on Text-to-Image (T2I) Generation with FLUX VAE (with and without CFG). We report metrics on GenEval (Ghosh et al., 2023) for 2B-parameter T2I models trained with our tokenizer and FLUX VAE. Each T2I model is trained for 100K steps and evaluated at $\times256$ resolution. Our tokenizer is trained on LAION dataset. The rFID is computed using 200K randomly sampled images from the COYO-700M dataset (Minwoo et al., 2022).
Tokenizer rFID Overall Single object Two object Counting Colors Position Color attribution
with CFG
FLUX VAE 0.102 0.476 0.978 0.477 0.346 0.813 0.107 0.135
Ours 0.443 0.556 0.984 0.702 0.446 0.824 0.120 0.257
without CFG
FLUX VAE 0.102 0.230 0.746 0.101 0.090 0.406 0.020 0.015
Ours 0.443 0.329 0.850 0.303 0.234 0.508 0.037 0.040
Table 14: Quantitative Comparison on Text-to-Image (T2I) Generation with VA-VAE (with and without CFG). We report metrics on diverse prompt sets for 1B-parameter T2I models trained with our tokenizer and the VA-VAE. Each T2I model is trained for 50K steps and evaluated at $\times256$ resolution. Both our tokenizer and VA-VAE are trained on ImageNet. The rFID is computed using 200K randomly sampled images from the COYO-700M dataset (Minwoo et al., 2022).
Coco Prompt 6K (Lin et al., 2014) (with CFG)
VA-VAE 0.241 34.13 0.016 0.221 0.426 -0.129 5.088 31.50 0.740
Ours 0.191 31.19 0.015 0.230 0.574 0.088 5.215 31.68 0.781
Parti Prompt (Yu et al., 2022) (with CFG)
VA-VAE 0.241 - - 0.220 0.444 -0.146 4.893 31.40 0.677
Ours 0.191 - - 0.228 0.556 0.062 5.015 31.73 0.706
HPSv2 Prompt (Wu et al., 2023) (with CFG)
VA-VAE 0.241 - - 0.201 0.435 -0.287 5.042 31.63 0.708
Ours 0.191 - - 0.210 0.565 -0.131 5.120 32.05 0.737
Coco Prompt 6K (Lin et al., 2014) (without CFG)
VA-VAE 0.241 42.35 0.019 0.179 0.448 -1.292 4.352 26.89 0.497
Ours 0.191 39.86 0.017 0.187 0.552 -1.121 4.493 27.53 0.548
Parti Prompt (Yu et al., 2022) (without CFG)
VA-VAE 0.241 - - 0.179 0.457 -1.396 4.163 26.23 0.493
Ours 0.191 - - 0.186 0.543 -1.213 4.320 26.99 0.524
HPSv2 Prompt (Wu et al., 2023) (without CFG)
VA-VAE 0.241 - - 0.162 0.455 -1.378 4.256 26.05 0.521
Ours 0.191 - - 0.170 0.545 -1.251 4.451 26.48 0.551
Table 15: GenEval Quanlitative Comparison on Text-to-Image (T2I) Generation with VA-VAE (with and without CFG). We report metrics on GenEval (Ghosh et al., 2023) for 1B-parameter T2I models trained with our tokenizer and VA-VAE. Each T2I model is trained for 50K steps and evaluated at $\times256$ resolution. Both our tokenizer and VA-VAE are trained on ImageNet. The rFID is computed using 200K randomly sampled images from the COYO-700M dataset (Minwoo et al., 2022).
Tokenizer rFID Overall Single object Two object Counting Colors Position Color attribution
with CFG
VA-VAE 0.241 0.411 0.944 0.338 0.275 0.782 0.050 0.078
Ours 0.191 0.454 0.978 0.434 0.338 0.795 0.073 0.110
without CFG
VA-VAE 0.241 0.194 0.653 0.071 0.088 0.335 0.010 0.005
Ours 0.191 0.243 0.734 0.136 0.125 0.418 0.033 0.013
C.5 More Qualitative Results
Fig. 17 and Fig. 18 show the comparison with FLUX-VAE (withou CFG) and VA-VAE (with CFG).